OpenStudy (anonymous):

One dimension of a rectangular map is 8 inches and the perimeter is 40 inches. What is the other dimension of the map? A. 14 inches B. 12 inches C. 4 inches D. 2 inches

OpenStudy (btaylor):

For a rectangle, the perimeter can be found with the formula P = 2W + 2L. If we let W=8 (arbitrarily, but it doesn't even matter), then we can solve for L.

OpenStudy (anonymous):

so would the answer be D?

OpenStudy (btaylor):

40 = 2(8) + 2L Solve for L.

OpenStudy (nirmalnema):

L=12 inches

OpenStudy (nirmalnema):

2(8+L)=40 8+L=20 L=20-8=12